OpenStudy (anonymous):

collect like terms 14u+5v-6u-7v=

OpenStudy (anonymous):

8u-2v

OpenStudy (bahrom7893):

14u-6u+5v-7v= 8u - 2v

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

is that supposed to be a - or a + in 8u 2v? why did it change from + to - in the answer

OpenStudy (anonymous):

5v-7v=-2v

OpenStudy (anonymous):

By like terms, they just mean sum the coefficients on each variable.. \[14u - 6u = (14 - 6)u = 8u\] etc.